jhonyy9 (jhonyy9):

true or false this proof ? - let p and k prims p=2a+1 k=2b+1 - a and b are number naturals so than a=(p-1)/2 and b=(k-1)/2 - n is number natural greater or equal 2 - prove that for every n there are a and b such that n=a+b+1 is true . n=a+b+1 a=(p-1)/2 b=(k-1)/2 so than n=(p-1)/2 +(k-1)/2 +1 is true - so n =(2a+1-1)/2 +(2b+1-1)/2 +1 n=2a/2 +2b/2 +1 n=a+b+1

jhonyy9 (jhonyy9):

- Be prime numbers p and t as P = (2x +1) and t = (2Y +1), where x and y are integers, - To show that whatever n, greater than or equal to 2, there are numbers x and y as x = (p-1) / 2 and y = (t-1) / 2, so that the equation n = x + y +1 is always true - For n = 2 we have p = 2 and t = 2 so that 2 = (2-1) / 2 + (2-1) / 2 +1 ie 2 = 1/2 +1 / 2 +1 or 2 = 1 + 1 or 2 = 2 - For n = 3 we have p = 3 and t = 3 so that 3 = (3-1) / 2 + (3-1) / 2 +1 or 3 = 1 + 1 +1 ie 3 = 3 - For. n = 4 we have p = 5 and t = 3 so that 4 = (5-1) / 2 + (3-1) / 2 +1 or 4 = 2 + 1 + 1 ie 4 = 4 ... - For. n = k we have k = x + y + 1 ie k = (p-1) / 2 + (t-1) / 2 +1 suppose is always true, so - For. k = k + 1 we have k +1 = ((p-1) / 2 + (t-1) / 2 +1) +1 ie k + 1 = k + 1 so q.e.d. - so can being accepted right,correct this proof ?